a partical moves in a straight line and its position x at time t is given by X^2=2+t. its acceleration given by...........?
 

Grade:11

1 Answers

Vikas TU
7 years ago
write X^2=2+t
as:
x = root(t +2)
differntiate it  to get velocity with respect to time.
dx/dt = v = 1/2root(t +2)
acceleration => a = dv/dt = – 1/4(t+2)^-3/2 at time t is the accln.